How many committees of two or more can be selected from 10 people?
Answer by richard1234(7193)

The number of ways is 10C2 + 10C3 + ... + 10C10. If you know that 10C0 + 10C1 + 10C2 + ... + 10C10 = 2^10 = 1024, we can take this and subtract off 10C0 and 10C1, leaving 1024 - 1 - 10 = 1013.